OpenStudy (anonymous):

Indefinite Integrals related problem: The acceleration due to gravity on Mars is approximately 3.72 m/sec2. If an astronaut jumps straight up on the surface of the planet with an initial velocity of 2 m/sec, what height will she attain? Find the comparable height that she would jump on the earth.

OpenStudy (jhannybean):

Can you use physics to solve this?

Miracrown (miracrown):

Perhaps

OpenStudy (jhannybean):

\[V_{f_y}^2 = V_{i_y}^2 + 2a_{y}\Delta y\]

OpenStudy (anonymous):

No. I've to use calculus.

Miracrown (miracrown):

We are given the acceleration due to gravity on Mars, which is approximately (3.72 m per second) per second. This is a constant, acceleration function.

Miracrown (miracrown):

Do you know how to find a velocity function on Mars ? @Erroneous_Copper

OpenStudy (jhannybean):

Oh, then possibly \[\int a(t)dt =v(t) +c \implies \int (v(t) +c)dt=s(t) +ct +d\] something like that perhaps.

OpenStudy (jhannybean):

and s(t) would be the position function (or height) at a given time.

Miracrown (miracrown):

v(t) , yes :) next, what is the antiderivative of 3.72 dt ?

Miracrown (miracrown):